# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ 4.2.2 Bounds for Residual Networks (ResNets)
$\begin{split}\alpha^{(l)}(\mathbf{x})&=\alpha^{(l-1)}(\mathbf{x})+\phi\left(\frac{1}{L\sqrt{m}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)~{},\\ f(\theta;\mathbf{x})&=\alpha^{(L+1)}(\mathbf{x})=\frac{1}{\sqrt{m}}\mathbf{v}^{\top}\alpha^{(L)}(\mathbf{x})~{},\end{split}$ (28)
where $\alpha^{(0)}(\mathbf{x})=\mathbf{x}$, and $W^{(l)},l\in[L],\mathbf{v},\phi$ are as in FFNs. Let $h^{(L)}(W,\mathbf{x}):=\frac{1}{\sqrt{m}}\alpha^{(L)}(\mathbf{x})$ be the “featurizer,” i.e., output of the last layer, so that $f(\theta;\mathbf{x})=\mathbf{v}^{\top}h^{(L)}(W,\mathbf{x})$ with $\theta=(W,\mathbf{v})$. Note that the scaling of the residual layers is assumed to have a smaller scaling factor $\frac{1}{L}$ following standard theoretical analysis of ResNets [21] and also aligns with practice where smaller initialization variance yields state-of-the-art performance [91]. With the featurizer and associated sets defined as in (26) for FFNs, we now establish an upper bound on the LGGW of gradients in terms of that of the featurizer.
Theorem 6 (Gaussian Width of Gradients: ResNets).
Under Assumptions 4 and 5, with $\beta:=\sigma_{1}+\frac{\rho}{\sqrt{m}}$, with probability at least $\left(1-\frac{2L}{m}\right)$ over the randomness of the initialization, we have
$w(\hat{\Xi}_{n}^{\operatorname{rn}})\leq c_{1}w(\hat{A}_{n}^{(L)})+c_{2}(1+\rho_{1})e^{\beta}~{}.$ (29)
@@
As for FFNs, choosing $\sigma_{1}<1-\frac{\rho}{\sqrt{m}}$, i.e., mildly small initialization variance, satisfies $\beta<1$. When $\beta<1$, we have $w(\hat{\Xi}_{n}^{\operatorname{rn}})=O(w(\hat{A}_{n}^{(L)}))$ since $(1+\rho_{1})e^{\beta}=O(1)$. Thus, for ResNets, the Gaussian width of gradients on all parameters reduces to the Gaussian width of the featurizer. ∎
@@
In Figure 2, we plot the average and max (over 5 runs) of the $L_{0}$ and $L_{1}$ norm of the featurizer for the last $10\%$ epochs close to convergence for three standard deep learning models: ResNet18 on CIFAR-10 (512 dimension featurizer), FFN on CIFAR-10 (256 dimension featurizer), and CNN on Fashion-MNIST (128 dimension featurizer). The small $L_{0}$ and $L_{1}$ norm of the featurizer, indicating both feature selection and being in small $L_{1}$ ball, imply small Gaussian width of the featurizer which in turn implies small LGGW based on Theorems 5 and 6. ∎
@@ 5 Conclusions
In this work, we have introduced a new approach to analyzing the generalization and optimization of learning models based on the loss gradient Gaussian width (LGGW). For loss satisfying the gradient domination condition, which includes the popular PL condition as a special case, we establish generalization bounds in terms of LGGW. In the context of optimization, we show that stochastic mini-batch gradient descent with sample reuse does not go haywire and the gradient estimates stay close to the population counterparts as long as the Gaussian width of the gradient is small. We also discuss the LGGW of deep learning models through empirical evidence, examples, and analysis of some special cases. We show that the LGGW can be bounded by the $L_{2}$ norm of the Hessian eigenvalues and the Gaussian width of the featurizer. Overall, our work demonstrates the benefits of small Gaussian width in the context of learning, and will hopefully serve as motivation for further study of this geometric perspective especially for modern learning models.
@@ Appendix A Generalization Bounds: Proofs for Section 2
In this section, we provide proofs for results in Section 2.
A.1 Generalization bound in terms NERC, following [27]
Proof sketch of Main Result
@@ A.2 Bounding NERC with Gaussian width
Then, an application of the bounded difference inequality (33[15][Theorem 6.2] completes the proof. ∎
@@
For the first step, note that there is a correspondence between $\hat{\xi}^{(n)}\in\hat{\Xi}^{(n)}$ and $\hat{\xi}_{n}\in\hat{\Xi}_{n}$ as $\hat{\xi}_{n}=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$, i.e., average of the $n$ gradients in $\hat{\xi}^{(n)}$. Moreover, for any pair $\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2}\in\hat{\Xi}^{(n)}$ respectively corresponding to $\hat{\xi}_{1,n},\hat{\xi}_{2,n}\in\hat{\Xi}_{n}$, by definition of $\bar{d}^{(n)}$ as in (17) and $d$ as in (14), we have
$\bar{d}^{(n)}(\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2})=d(\hat{\xi}_{1,n},\hat{\xi}_{2,n})~{}.$ (42)
@@
For the second step, consider the Gaussian process $X_{\hat{\xi}_{n}}=\langle\hat{\xi}_{n},\mathbf{g}\rangle$ on $\hat{\Xi}_{n}$. The canonical distance for the Gaussian process is
@@
which is exactly $d$ as in (14). As a result, from the majorizing measure theorem [80], for some constant $c_{4}$, we have
@@ B.1 GD with sample reuse
$\Lambda_{t}\triangleq\left\{\sup_{\theta_{t}\in\Theta}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}\leq\epsilon\right\}~{},~{}~{}\text{and}~{}~{}\bar{\Lambda}_{t}\triangleq\left\{\sup_{\theta_{t}\in\Theta}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}>\epsilon\right\}~{}.$ (48)
@@
Note that (posterior) probability of any event $\lambda$ at stage $1$, i.e., conditioned on $\Lambda_{1}$ of the filtration can be written in terms of probabilities at the previous stage based on the definition of conditional probability as
$\mathbb{P}_{{\cal D}_{1}}[\lambda]=\mathbb{P}_{{\cal D}|\{\Lambda_{1}\}}[\lambda]=\frac{\mathbb{P}_{{\cal D}}[\lambda,\Lambda_{1}]}{\mathbb{P}_{{\cal D}}[\Lambda_{1}]}~{}.$ (51)
@@
$\mathbb{P}_{{\cal D}_{t}}[\lambda]=\mathbb{P}_{{\cal D}|\{\Lambda_{\tau\leq t}\}}[\lambda]=\frac{\mathbb{P}_{{\cal D}}[\lambda,\{\Lambda_{\tau\leq t}\}]}{\mathbb{P}_{{\cal D}}[\{\Lambda_{\tau\leq t}\}]}~{},$ (52)
where $\{\Lambda_{\tau\leq t}\}$ denotes the joint event $\{\Lambda_{1},\ldots,\Lambda_{t}\}$ where individual events $\Lambda_{\tau},\tau\in[t]$ differ based on $\theta_{\tau}$ for that event, and $\mathbb{P}_{{\cal D}}[\{\Lambda_{\tau\leq t}\}]$ is the probability of the joint according to the original distribution ${\cal D}$. While seemingly straightforward, (54) shows a way to express probabilities $\mathbb{P}_{{\cal D}_{t}}[\cdot]$ of events in the stochastic process directly in terms $\mathbb{P}_{{\cal D}}[\cdot]$, the original distribution.
For our analysis, instead of working with the full distributions, we will be working on the finite sample versions based on $z^{(n)}\sim{\cal D}^{n}$. In particular, let ${\cal D}_{0}(z^{(n)})$ be the empirical distribution based on the samples, i.e., ${\cal D}_{0}(z^{(n)})=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}_{z_{i}}$, where $\mathbbm{1}$ is the indicator function. As before, the (posterior) probability of any event $\lambda$ at any stage $1$ of the filtration can be written in terms of probabilities at the previous stage based on the definition of conditional probability:
$\mathbb{P}_{{\cal D}_{1}(z^{(n)})}[\lambda]=\mathbb{P}_{{\cal D}_{0}(z^{(n)})|\{\Lambda_{1}\}}[\lambda]=\frac{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\lambda,\Lambda_{1}]}{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\Lambda_{1}]}~{}.$ (53)
@@
$\mathbb{P}_{{\cal D}_{t}(z^{(n)})}[\lambda]=\mathbb{P}_{{\cal D}_{0}(z^{(n)})|\{\Lambda_{\tau\leq t}\}}[\lambda]=\frac{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\lambda,\{\Lambda_{\tau\leq t}\}]}{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\{\Lambda_{\tau\leq t}\}]}~{}.$ (54)
@@
The focus of our work is to bound to following probability
To simplify the notation, let random variable ${X}_{t}\triangleq\sup_{\theta_{t}\in\Theta}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}$. We have events $\Lambda_{t}=\{X_{t}\leq\epsilon\}$ and $\bar{\Lambda}_{t}=\{X_{t}>\epsilon\}$, for $t\in[T]$. Thus we have
@@
$\mathbb{P}_{{\cal D}_{T}(z^{(n)})}\left[\max_{t\in[T]}\sup_{\theta_{t}\in\Theta}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}>\epsilon\right]$ $\leq\sum_{t=1}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}\left[\sup_{\theta_{t}\in\Theta}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}>\epsilon\right]$
$=\sum_{t=1}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}\left[X_{t}>\epsilon\right]$
$=\sum_{t=1}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}[\bar{\Lambda}_{t}]$ (55)
The unique aspect in developing such a large deviation bound is the need to acknowledge and incorporate the fact that the sequence $\theta_{1:T}$ was computed sequentially, i.e., any $\theta_{t}$ depends on $\{\theta_{\tau<t}\}$. We have
@@
$\mathbb{P}_{{\cal D}_{T}(z^{(n)})}$ $\left[\max_{t\in[T]}\sup_{\theta_{t}\in\Theta}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}>\epsilon\right]$
$=\sum_{t=1}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}[\bar{\Lambda}_{t}]$
$=\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\bar{\Lambda}_{1}]+\mathbb{P}_{{\cal D}_{1}(z^{(n)})}[\bar{\Lambda}_{2}]+\ldots+\mathbb{P}_{{\cal D}_{T-1}(z^{(n)})}[\bar{\Lambda}_{T}]$
$=\mathbb{P}_{z^{(n)}}[\bar{\Lambda}_{1}]+\frac{\mathbb{P}_{z^{(n)}}[\Lambda_{1},\bar{\Lambda}_{2}]}{\mathbb{P}_{z^{(n)}}[\Lambda_{1}]}+\ldots+\frac{\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\},\bar{\Lambda}_{T}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\}]}$
$\overset{(a)}{\leq}\frac{\mathbb{P}_{z^{(n)}}[\bar{\Lambda}_{1}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\}]}+\frac{\mathbb{P}_{z^{(n)}}[\Lambda_{1},\bar{\Lambda}_{2}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\}]}+\ldots+\frac{\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\},\bar{\Lambda}_{T}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\}]}$
$\overset{(b)}{\leq}\frac{1-\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq T}\}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\}]}$
$=\frac{1-\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}\leq\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T-1]}X_{t}\leq\epsilon\right]}$
$=\frac{\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}>\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T-1]}X_{t}\leq\epsilon\right]}$
$\overset{(c)}{\leq}\frac{\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}>\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}\leq\epsilon\right]}~{}.$ (56)
@@
where (a) follows since $\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq(T-1)}\}]\leq\mathbb{P}_{z^{(n)}}[\{\Lambda_{t\leq\tau}\}]\leq 1$ for all $\tau\in[T]$, (b) follows since the counter-event of $\{\Lambda_{t\leq T}\}$ covers the events in $(a)$, and (c) follows since $\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T]}X_{t}\leq\epsilon\right]\leq\mathbb{P}_{z^{(n)}}\left[\max_{t\in[T-1]}X_{t}\leq\epsilon\right]$. Thus, it suffices to focus on an upper bound on $\mathbb{P}_{z^{(n)}\sim{\cal D}^{n}}[\max_{t\in[T]}X_{t}>\epsilon]$ which corresponds to the same ‘bad’ event $\{\max_{t\in[T]}X_{t}>\epsilon\}$ but in the non-adaptive setting, since the probability is w.r.t. $\mathbb{P}_{z^{(n)}\sim{\cal D}^{n}}\left[\cdot\right]$. Now, note that

### v2

@@ 4.2.2 Bounds for Residual Networks (ResNets)
$\begin{split}\alpha^{(l)}(\mathbf{x})&=\alpha^{(l-1)}(\mathbf{x})+\phi\left(\frac{1}{L\sqrt{m_{l}}}W^{(l)}\alpha^{(l-1)}(\mathbf{x})\right)~{},\\ f(\theta;\mathbf{x})&=\alpha^{(L+1)}(\mathbf{x})=\mathbf{v}^{\top}\alpha^{(L)}(\mathbf{x})~{},\end{split}$ (29)
where $\alpha^{(0)}(\mathbf{x})=\mathbf{x}$, and $W^{(l)},l\in[L],\mathbf{v},\phi$ are as in FFNs. Let $h^{(L)}(W,\mathbf{x}):=\alpha^{(L)}(\mathbf{x})\in\mathbb{R}^{m_{L}}$ be the “featurizer,” i.e., output of the last layer, so that $f(\theta;\mathbf{x})=\mathbf{v}^{\top}h^{(L)}(W,\mathbf{x})$ with $\theta=(W,\mathbf{v})$. Note that the scaling of the residual layers is assumed to have a smaller scaling factor $\frac{1}{L}$ following standard theoretical analysis of ResNets [22] and also aligns with practice where smaller initialization variance yields state-of-the-art performance [92]. With the featurizer and associated sets defined as in (27) for FFNs, we now establish an upper bound on the single-sample LGGW of ResNet gradients in terms of that of the featurizer.
Theorem 5 (LGGW: ResNets).
Under Assumptions 3 and 4, with $\beta_{l}:=\sigma_{1}+\frac{\rho}{\sqrt{m_{l}}}$, $l\in[L]$, with probability at least $\left(1-\sum_{l=1}^{L}\frac{2}{m_{l}}\right)$ over the randomness of the initialization, we have
$w(\Xi^{\operatorname{rn}})\leq c_{1}w(A^{(L)})+c_{2}\frac{1+\rho_{1}}{L}\sqrt{m_{L}}\prod_{l=1}^{L}\left(1+\frac{\beta_{l}}{L}\right)\sum_{l=1}^{L}\frac{1}{\left(1+\frac{\beta_{l}}{L}\right)\sqrt{m_{l}}}.$ (30)
@@
For the simplest case, with $m_{l}=m$ for all $l\in[L]$, we have that $w(\Xi^{\operatorname{rn}})\leq c_{1}w(A^{(L)})+c_{2}\left(1+\rho_{1}\right)\left(1+\frac{\beta}{L}\right)^{L-1}\leq c_{1}w(A^{(L)})+c_{2}(1+\rho_{1})e^{\beta}$. in which $\beta=\sigma_{1}+\frac{\rho}{\sqrt{m}}$, and the second inequality holds since $\left(1+\frac{\beta}{L}\right)^{L-1}\leq e^{\beta}$. As for FFNs, choosing $\sigma_{1}<1-\frac{\rho}{\sqrt{m}}$, i.e., mildly small initialization variance, satisfies $\beta<1$. When $\beta<1$, we have $w(\Xi^{\operatorname{rn}})=O(w(A^{(L)}))$ since $(1+\rho_{1})e^{\beta}=O(1)$. Thus, for ResNets, the LGGW on all parameters reduces to the Gaussian width of the featurizer. ∎
@@
As for FFNs, our ResNet bound is also much sharper when the dimensions in the hidden layers are much larger than the output dimension, which is the case in commonly used ResNet models. ∎
Remark 4.6.
To get a sense of the Gaussian width of the featurizer, which will be small if the featurizer is sparse [84], in Figure 3, we plot the average and max (over 5 runs) of the $L_{0}$ and $L_{1}$ norm of the featurizer for the last $10\%$ epochs close to convergence for three standard deep learning models: ResNet18 on CIFAR-10 (512 dimension featurizer), FFN on CIFAR-10 (256 dimension featurizer), and CNN on Fashion-MNIST (128 dimension featurizer). The small $L_{0}$ and $L_{1}$ norm of the featurizer, indicating both feature selection and being in small $L_{1}$ ball, imply small Gaussian width of the featurizer which in turn implies small LGGW based on Theorems 4 and 5. ∎
@@ 5 Conclusions
In this work, we have introduced a new approach to analyzing the generalization and optimization of learning models based on the loss gradient Gaussian width (LGGW). For loss satisfying the gradient domination condition, which includes the popular PL condition as a special case, we establish generalization bounds in terms of LGGW. In the context of optimization, we show that gradient descent with sample reuse does not go haywire and the gradient estimates stay close to the population counterparts as long as the LGGW at initialization is small. We also consider the LGGW of deep learning models, especially feedforward networks and residual networks. In these settings, we show that the LGGW with single sample can be bounded by the Gaussian width of the featurizer and additional terms which are small under mild conditions. Overall, our work demonstrates the benefits of small loss gradient Gaussian width (LGGW) in the context of learning, and will hopefully serve as motivation for further study of this geometric perspective especially for modern learning models.
@@ Appendix A Generalization Bounds: Proofs for Section 2
In this section, we first provide a proof sketch of the main results. Then we provide detail proofs for results in Section 2.
A.1 Generalization bound in terms of NERC
@@ A.2 Bounding NERC with Gaussian width
Then, an application of the bounded difference inequality (34[16][Theorem 6.2] completes the proof. ∎
@@
For the first step, note that there is a correspondence between $\hat{\xi}^{(n)}\in\hat{\Xi}^{(n)}$ and $\hat{\xi}_{n}\in\hat{\Xi}_{n}$ as $\hat{\xi}_{n}$ is a scaled stacked version of the $n$-tuples. Moreover, for any pair $\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2}\in\hat{\Xi}^{(n)}$ respectively corresponding to $\hat{\xi}_{1,n},\hat{\xi}_{2,n}\in\hat{\Xi}_{n}$, by definition of $\bar{d}^{(n)}$ as in (17) and $d$ as in (14), we have
$d(\hat{\xi}_{1,n},\hat{\xi}_{2,n})$ $=\mathbb{E}\left[\left(X_{\hat{\xi}_{1,n}}-X_{\hat{\xi}_{2,n}}\right)^{2}\right]$
$=\frac{1}{n}\mathbb{E}\left[\left(\langle\hat{\xi}_{1,n},g\rangle-\langle\hat{\xi}_{2,n},g\rangle\right)^{2}\right]$
$=\frac{1}{n}\mathbb{E}\left[\langle\hat{\xi}_{1,n}-\hat{\xi}_{2,n},g\rangle^{2}\right]$
$=\frac{1}{n}\sum_{i=1}^{n}\|\xi_{i,1}-\xi_{i,2}\|_{2}^{2}$
$=\bar{d}^{(n)}(\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2})~{}.$
@@
For the second step, since for the Gaussian process $X_{\hat{\xi}_{n}}=\langle\hat{\xi}_{n},\mathbf{g}\rangle$ on $\hat{\Xi}_{n}$, the canonical distance is
@@
from the majorizing measure theorem [81], for some constant $c_{4}$, we have
@@ B.1 GD with sample reuse
$\Lambda_{0}\triangleq\left\{\sup_{\theta_{0}\in\Theta_{0}}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{0},z_{i})-\nabla{\cal L}_{D}(\theta_{0})\bigg{\|}_{2}\leq\epsilon\right\}~{},$ $~{}~{}\text{and}~{}~{}\bar{\Lambda}_{0}\triangleq\left\{\sup_{\theta_{0}\in\Theta_{0}}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{0},z_{i})-\nabla{\cal L}_{D}(\theta_{0})\bigg{\|}_{2}>\epsilon\right\}~{},$ (48)
$\Lambda_{t}\triangleq\left\{\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}\leq\epsilon\right\}~{},$ $~{}~{}\text{and}~{}~{}\bar{\Lambda}_{t}\triangleq\left\{\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}>\epsilon\right\}~{}.$ (49)
@@
Note that (posterior) probability of any event $\lambda_{1}$ at stage $1$, i.e., conditioned on $\Lambda_{1}$ of the filtration can be written in terms of probabilities at the previous stage based on the definition of conditional probability as
$\mathbb{P}_{{\cal D}_{1}}[\lambda_{1}]=\mathbb{P}_{{\cal D}|\{\Lambda_{0}\}}[\lambda_{1}]=\frac{\mathbb{P}_{{\cal D}}[\lambda_{1},\Lambda_{0}]}{\mathbb{P}_{{\cal D}}[\Lambda_{0}]}~{}.$ (52)
@@
$\mathbb{P}_{{\cal D}_{t}}[\lambda]=\mathbb{P}_{{\cal D}|\{\Lambda_{\tau<t}\}}[\lambda_{t}]=\frac{\mathbb{P}_{{\cal D}}[\lambda_{t},\{\Lambda_{\tau<t}\}]}{\mathbb{P}_{{\cal D}}[\{\Lambda_{\tau<t}\}]}~{},$ (53)
where $\{\Lambda_{\tau<t}\}$ denotes the joint event $\{\Lambda_{0},\Lambda_{1},\ldots,\Lambda_{t-1}\}$ where individual events $\Lambda_{\tau}$ are as in (48) and (49). While seemingly straightforward, (53) shows a way to express probabilities $\mathbb{P}_{{\cal D}_{t}}[\cdot]$ of events in the sequence directly in terms $\mathbb{P}_{{\cal D}}[\cdot]$, the original distribution.
For our analysis, instead of working with the full distributions, we will be working on the finite sample versions based on $z^{(n)}\sim{\cal D}^{n}$. In particular, let ${\cal D}_{0}(z^{(n)})$ be the empirical distribution based on the samples, i.e., ${\cal D}_{0}(z^{(n)})=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}_{z_{i}}$, where $\mathbbm{1}$ is the indicator function. As before, the (posterior) probability of any event $\lambda_{1}$ at any stage $1$ of the filtration can be written in terms of probabilities at the previous stage based on the definition of conditional probability:
$\mathbb{P}_{{\cal D}_{1}(z^{(n)})}[\lambda_{1}]=\mathbb{P}_{{\cal D}_{0}(z^{(n)})|\{\Lambda_{0}\}}[\lambda_{1}]=\frac{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\lambda_{1},\Lambda_{0}]}{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\Lambda_{0}]}~{}.$ (54)
@@
$\mathbb{P}_{{\cal D}_{t}(z^{(n)})}[\lambda_{t}]=\mathbb{P}_{{\cal D}_{0}(z^{(n)})|\{\Lambda_{\tau\leq t}\}}[\lambda_{t}]=\frac{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\lambda_{t},\{\Lambda_{\tau<t}\}]}{\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\{\Lambda_{\tau<t}\}]}~{}.$ (55)
@@
To simplify the notation, consider the random variable
${X}_{0}$ $\triangleq\sup_{\theta_{0}\in\Theta_{0}}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{0},z_{i})-\nabla{\cal L}_{D}(\theta_{0})\bigg{\|}_{2}~{},$ (56)
${X}_{t}$ $\triangleq\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}~{}.$ (57)
Then, from (48) and (49), we have $\Lambda_{t}=\{X_{t}\leq\epsilon\}$ and $\bar{\Lambda}_{t}=\{X_{t}>\epsilon\}$. Then, we have
@@
$\mathbb{P}$ $\left[\max\left(\sup_{\theta_{0}\in\Theta_{0}}\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{0},z_{i})-\nabla{\cal L}_{D}(\theta_{0})\right\|_{2}~{},~{}~{}~{}\max_{t\in[T]}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}\right)>\epsilon\right]$ (58)
@@
$\leq\mathbb{P}_{z^{(n)}}\left[\sup_{\theta_{0}\in\Theta_{0}}\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{0},z_{i})-\nabla{\cal L}_{D}(\theta_{0})\bigg{\|}_{2}>\epsilon\right]+\sum_{t=1}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}\left[\bigg{\|}\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\bigg{\|}_{2}>\epsilon\right]$
$=\sum_{t=0}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}\left[X_{t}>\epsilon\right]$
$=\sum_{t=0}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}[\bar{\Lambda}_{t}]~{}.$ (59)
Now, with $\{\Lambda_{0:\tau}\}:=(\Lambda_{0},\Lambda_{1},\ldots,\Lambda_{\tau})$ and $\{0:\tau\}:=\{0,1,\ldots,\tau\}$, we have
$\sum_{t=0}^{T}\mathbb{P}_{{\cal D}_{t}(z^{(n)})}[\bar{\Lambda}_{t}]$ $=\mathbb{P}_{{\cal D}_{0}(z^{(n)})}[\bar{\Lambda}_{0}]+\mathbb{P}_{{\cal D}_{1}(z^{(n)})}[\bar{\Lambda}_{1}]+\ldots+\mathbb{P}_{{\cal D}_{T}(z^{(n)})}[\bar{\Lambda}_{T}]$
$=\mathbb{P}_{z^{(n)}}[\bar{\Lambda}_{0}]+\frac{\mathbb{P}_{z^{(n)}}[\Lambda_{0},\bar{\Lambda}_{1}]}{\mathbb{P}_{z^{(n)}}[\Lambda_{0}]}+\ldots+\frac{\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\},\bar{\Lambda}_{T}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\}]}$
$\overset{(a)}{\leq}\frac{\mathbb{P}_{z^{(n)}}[\bar{\Lambda}_{0}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\}]}+\frac{\mathbb{P}_{z^{(n)}}[\Lambda_{0},\bar{\Lambda}_{1}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\}]}+\ldots+\frac{\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\},\bar{\Lambda}_{T}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\}]}$
$\overset{(b)}{\leq}\frac{1-\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:T}\}]}{\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\}]}$
$=\frac{1-\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T\}}X_{t}\leq\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T-1\}}X_{t}\leq\epsilon\right]}$
$=\frac{\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T\}}X_{t}>\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T-1\}}X_{t}\leq\epsilon\right]}$
$\overset{(c)}{\leq}\frac{\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T\}}X_{t}>\epsilon\right]}{\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T\}}X_{t}\leq\epsilon\right]}~{}.$ (60)
@@
where (a) follows since $\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:(T-1)}\}]\leq\mathbb{P}_{z^{(n)}}[\{\Lambda_{0:\tau}\}]\leq 1$ for all $\tau\in\{0:T\}$, (b) follows since the counter-event of $\{\Lambda_{0:T}\}$ covers the events in $(a)$, and (c) follows since $\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T\}}X_{t}\leq\epsilon\right]\leq\mathbb{P}_{z^{(n)}}\left[\max_{t\in\{0:T-1\}}X_{t}\leq\epsilon\right]$. Thus, it suffices to focus on an upper bound on $\mathbb{P}_{z^{(n)}\sim{\cal D}^{n}}[\max_{t\in\{0:T\}}X_{t}>\epsilon]$ which corresponds to the same ‘bad’ event $\{\max_{t\in\{0:T\}}X_{t}>\epsilon\}$ but in the non-adaptive setting, since the probability is w.r.t. $\mathbb{P}_{z^{(n)}\sim{\cal D}^{n}}\left[\cdot\right]$.
From Lemma 3, for any $u>0$, we have